OpenStudy (anonymous):

Solve: 6x − 10 + 14 = 22

Parth (parthkohli):

Operate on the like terms. \(\Large \color{MidnightBlue}{\Rightarrow 6x + 4 = 22 }\) Subtract 4 from both sides. \(\Large \color{MidnightBlue}{\Rightarrow 6x = 18 }\) Divide both sides by 6. \(\Large \color{MidnightBlue}{\Rightarrow x = 3 }\)

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

6x=22+10-14=18 x=18/6=3

OpenStudy (anonymous):

6x-10+14 = 22 6x-4+4= 22+4 6x = 26 6x/6 = 26/6 x = 4.3

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

@dahmanman sorry ur wrong

OpenStudy (anonymous):

yeah whoops